The present age of a man is 3 times that of his son. 6 years ago, the age of the man was 4 times that of his son. Find the ratio of their ages 6 years later.
Answer by mananth(16946)

The present age of a man is 3 times that of his son. 6 years ago, the age of the man was 4 times that of his son. Find the ratio of their ages 6 years later.
son = x
father = 3x
..
6 years ago
son = x-6
father = 3x-6
..
3x-6=4(x-6)
3x-6=4x-24
-x=-18
x=18 son's age
father's age = 54 years